Circular No. 8996

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

CBAT@CFA.HARVARD.EDU (science)

URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 T5 (NEAT)

K. Kadota, Ageo, Saitama-ken, Japan, reports his recovery of comet P/2001 J1 (cf. IAUCs 7623, 7625, 7635), showing a central condensation, no tail, and coma diameter 0′.5–0′.6 on CCD images obtained with a 0.25-m f/6 reflector. Confirming observations were also made by H. Abe (Yatsuka-cho, Shimane-ken, Japan, 0.26-m reflector).

2008 UT α_{2000} δ_{2000} Mag. Observer
Oct. 15.77691 9^{h}28^{m}55.^{s}24 +5^{o}24′59′′.0  Kadota
15.77815 9 28 55.63 +5 24 56.6  _{”}
15.77939 9 28 56.04 +5 24 54.2 16.4  _{”}
17.80022 9 39 16.67 +4 26 46.8  _{”}
17.80196 9 39 17.19 +4 26 44.0  _{”}
17.80365 9 39 17.68 +4 26 41.2 16.5  _{”}
18.79792 9 44 20.38 +3 58 02.9  Abe
18.80833 9 44 23.43 +3 57 45.7 16.0  _{”}
18.82345 9 44 28.09 +3 57 19.3  Kadota
18.82525 9 44 28.64 +3 57 15.9  _{”}
18.82704 9 44 29.20 +3 57 12.9 16.4  _{”}

The indicated correction to the prediction on MPC 54170 is ΔT = −0.6 day. The following orbital elements by B. G. Marsden are from 68 observations, 2000–2008 (mean residual 0′′.9):

Epoch = 2001 Apr. 1.0 TT

T = 2001 Mar. 14.1052 TT  ω = 271.^{o}0222
e = 0.758290  Ω = 200.7993 }2000.0
q = 0.937075 AU  i = 10.1577

a = 3.876861 AU n^{o} = 0.1291171 P = 7.63 years

Epoch = 2008 Oct. 21.0 TT

T = 2008 Nov. 6.2551 TT  ω = 271.^{o}1732
e = 0.757141  Ω = 200.6744 }2000.0
q = 0.944153 AU  i = 10.1500

a = 3.887655 AU n^{o} = 0.1285797 P = 7.66 years

V1251 CYGNI

M. Templeton, AAVSO, reports a rare outburst of this dwarf nova (cf. IAUC 6123); P. Schmeer reports visual mag 12.6 on Oct. 19.871 UT.

2008 October 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green